OpenStudy (anonymous):

90% of what number is 45? I got 50 is that correct

OpenStudy (anonymous):

yes it is 50*0.9=45

OpenStudy (anonymous):

45/0.9= 50 that is how I got my answer

OpenStudy (anonymous):

yes, that comes from x*0.9=45